OpenStudy (kendricklamar2014):

Simplify: 8P3 A. 42 B. 336 C. 40,432 D. 56

OpenStudy (shamim):

nPr=n!/(n-r)!

OpenStudy (shamim):

8P3=8!/(8-3)!

OpenStudy (yanasidlinskiy):

Do you understand how to do it? @KendrickLamar2014

OpenStudy (shamim):

5!=?

OpenStudy (yanasidlinskiy):

The formula you would use is \[nPr=\frac{ n! }{ (n-r)! }\]

OpenStudy (kendricklamar2014):

yes

OpenStudy (yanasidlinskiy):

Next you would \[\frac{ 8! }{ (8-3)! } = \frac{ 8! }{ 5! } = \frac{ 8*7*6*5*4*3*2*1 }{ 5*4*3*2*1 }\]

OpenStudy (anonymous):

That is a partial permutation, \[nPk=\frac{ n! }{ (n-k)! } \rightarrow 8P3 =\frac{ 8! }{ (8-3)! }=\frac{ 8! }{ 5! }= 6\times7\times8= 336\]

OpenStudy (kendricklamar2014):

So the answer is B

OpenStudy (anonymous):

Yes

OpenStudy (yanasidlinskiy):

DO you have a calculator handy? If so, Use of what I did. Multiply the top and multiply the bottom. Then divide it and that would be your final answer!:) Good Luck!

OpenStudy (kendricklamar2014):

40320/120= 336

OpenStudy (anonymous):

It's easier if you cancel them out

OpenStudy (anonymous):

5! = 5*4*3*2*1

OpenStudy (anonymous):

and 8! = 8*7*6*5*4*3*2*1

OpenStudy (anonymous):

Just cancel the 5*4*3*2*1 Because they`re dividing and it is multiplying both

OpenStudy (kendricklamar2014):

Ok, thanks

OpenStudy (shamim):

welcome